OpenStudy (anonymous):

x<-2 or x>4 write interval notation for each compound inequality

OpenStudy (anonymous):

First: \[ (-\infty, -2) \] Second: \[ (4,\infty) \]

jimthompson5910 (jim_thompson5910):

then combine into \[\Large \left(-\infty,-2\right)\cup\left(4,\infty\right)\] The U signifies "or"